The function f(x)= [x]- x is not integrable in [ 0,3] , where [ x] denote greatest integer function.





True or false with full explanation

1
Expert's answer
2022-02-25T07:49:38-0500

We remind that the function "[x]" has on the interval "[0,3]" the following form: "[x]=0," "0\leq x\leq1", "[x]=1,1\leq x\leq2", "[x]=2,2\leq x\leq3". It is clear that the function "f(x)=[x]-x" is integrable on intervals "[0,1],[1,2]" and "[2,3]". Thus the function "x-[x]" is integrable on the interval"[0,3]", because it is bounded and has finitely many points of discontinuity.